A buffer is prepared using 0.28 M HNO2 and 0.23 M NO2 − . The pKa for HNO2 is 3.34. What is the pH of

this buffer after the addition of 0.05 M HCl?


1
Expert's answer
2022-02-01T09:12:02-0500

"pH = pKa + log10([A-]\/[HA])"

= 3.34 + log (0.33/0.23)

= 3.34+ 0.2= 3.54